OpenStudy (anonymous):

Which of the following is the arithmetic sequence formula for this list? 16, 31, 46, 61, ... i think its the third one.........

OpenStudy (anonymous):

What is the difference between : The2nd term and the first? 31-16? The 3rd and the 2nd: 46-31? The 4th and the 3rd: 61-46?

OpenStudy (anonymous):

sorry i didnt post tghe answers i think its tn=n+16

OpenStudy (anonymous):

@philo1234

OpenStudy (anonymous):

What is the answer to my questions above. What did you get for the three calculations?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

31-16 =15

OpenStudy (anonymous):

haha oops..

OpenStudy (anonymous):

tn=15n+1??

OpenStudy (anonymous):

tn=n+15. Do u have that as one of the choices?

OpenStudy (anonymous):

the choices are tn=15n+1 and tn=16n+31 and tn=n+16

OpenStudy (anonymous):

@philo1234

OpenStudy (anonymous):

@AriPotta @jim_thompson5910

jimthompson5910 (jim_thompson5910):

when n = 1, tn should be 16 so let's check tn=16n+31 tn=16n+31 tn=16(1)+31 ... Plug in n = 1 tn = 47 which is NOT 16, so tn=16n+31 is NOT the answer

OpenStudy (anonymous):

so what is the answer??

jimthompson5910 (jim_thompson5910):

you need to check n = 1 and n = 2 for each answer choice and you should get 16 and 31

OpenStudy (anonymous):

Jim is correct I was approaching the question with the wrong view. 15n+1 would be the correct answer. if you do the same check that jim did for 16n+31. check when n=1, 2, 3 and so on.

OpenStudy (anonymous):

oh ok thanks:)